requency Tables Due Sun 02/03/2019 10:00am 70 adults with gum disease were asked the number of times per week they used to floss before their diagnoses. The (incomplete) results are shown below: # of times floss per week Frequency Relative Frequency 9 0.1286 Cumulative Frequency 0.1571 0.1 0.0143 28 0.1143 0.1429 0.1857 a. Complete the table (Use 4 decimal places when applicable) 13 70 b. What is the cumulative relative frequency for flossing 7 times per week? Hint: Frequency Tables Textbook Pages Points possible: 1 This ts attempt 1 of 3. Nessage instructor about this question MacBook A 8 0
Answer #1

Solution:-

a. Explanation

1 11 0.1571 20 (9+11) = 20

4 11 0.1571 39 11/70 = 0.1571

6 10 0.1429 57 70 - (47 + 13) = 10


b. The cummulative relative frequency for flossing & times per week is 100%

=> (70/70)*100 = 100%